Scientists would like to place an artificial satellite into orbit at a height of 23,025 miles directly above the Earth's equator. Recall that the radius of the Earth is approximately 4,000 miles. Use the formula for the circumference of a circle to find the length of the satellite's orbit.
ROUND the answer to the nearest WHOLE number.
I am really struggling with this question... Please help! Thank you!

Answer by lwsshak3(11628)

Scientists would like to place an artificial satellite into orbit at a height of 23,025 miles directly above the Earth's equator. Recall that the radius of the Earth is approximately 4,000 miles. Use the formula for the circumference of a circle to find the length of the satellite's orbit.
ROUND the answer to the nearest WHOLE number.
**
Circumference of a circle=2π*radius
For given problem:
radius=23025+4000=27025 miles
length of the satellite's orbit=2π*27025=169,803 miles